ABSORPTION LINE SURVEY OF H_{3}^{+} TOWARD THE GALACTIC CENTER

\underline{MIWA GOTO}, Subaru Telescope, 650, North A’ohoku Place, Hilo HI 96720/ Institute for Astronomy, University of Hawaii, 640 North A’ohoku Place, Hilo HI 96720; B. J. McCALL, Department of Chemistry and Department of Astronomy, University of California, 601, Campbell Hall, Berkeley, CA 94720; T. R. GEBALLE, Gemini Observatory, 670, North A’ohoku Place, Hilo HI 96720T. USUDASubaru Telescope, 650, North A’ohoku Place, Hilo HI 96720; TAKESHI OKA, Department of Chemistry and Department of Astronomy and Astrophysics, the Enrico Fermi Institute, the University of Chicago, Chicago, IL 60637.

We present high-resolution (R = 20000) spectroscopy of H_{3}^{+} absorption toward the luminous Galactic center sources GCS 3-2 and GC IRS 3. With the efficient wavelength coverage afforded by a new generation spectrograph available at the 8-m class telescope, six absorption lines of H_{3}^{+} have been detected in each source, three of which are new. In particular the 3.543 μm absorption line of the R(3, 3)^{l} transition arising from the metastable (J, K) = (3, 3) state has been tentatively detected for the first time in the interstellar medium. The H_{3}^{+} absorption toward the Galactic center takes place in dense and diffuse clouds along the line of sight as well as the molecular complex close to the Galactic nucleus. At least four kinematic components are found in the H_{3}^{+} absorption lines. We suggest identifications of the velocity components with those of H i, CO, and H_{2}CO previously reported from radio and infrared observations. Our observation has revealed a striking difference between the absorption profiles of H_{3}^{+} and CO, demonstrating that the spectroscopy of H_{3}^{+} provides information complementary to that obtained from CO spectroscopy.